# Frequency-Assisted Mamba for Remote Sensing Image Super-Resolution

### v1

@@ Abstract
Recent progress in remote sensing image (RSI) super-resolution (SR) has exhibited remarkable performance using deep neural networks, e.g., Convolutional Neural Networks and Transformers. However, existing SR methods often suffer from either a limited receptive field or quadratic computational overhead, resulting in sub-optimal global representation and unacceptable computational costs in large-scale RSI. To alleviate these issues, we develop the first attempt to integrate the Vision State Space Model (Mamba) for RSI-SR, which specializes in processing large-scale RSI by capturing long-range dependency with linear complexity. To achieve better SR reconstruction, building upon Mamba, we devise a Frequency-assisted Mamba framework, dubbed FMSR, to explore the spatial and frequent correlations. In particular, our FMSR features a multi-level fusion architecture equipped with the Frequency Selection Module (FSM), Vision State Space Module (VSSM), and Hybrid Gate Module (HGM) to grasp their merits for effective spatial-frequency fusion. Recognizing that global and local dependencies are complementary and both beneficial for SR, we further recalibrate these multi-level features for accurate feature fusion via learnable scaling adaptors. Extensive experiments on AID, DOTA, and DIOR benchmarks demonstrate that our FMSR outperforms state-of-the-art Transformer-based methods HAT-L in terms of PSNR by 0.11 dB on average, while consuming only 28.05% and 19.08% of its memory consumption and complexity, respectively.
@@ I Introduction
High-resolution remote sensing imagery (RSI), which records high-quality earth observation details, provides promising prospects for large-scale and fine-grained applications [1, 2, 3, 4, 5]. However, the complex imaging environment (e.g., scattering and tremor) often impedes high-resolution (HR) image acquisition [6, 7, 8]. Moreover, RSI often undergoes severe compression and downsampling to tame the transmission instability between satellites and ground stations, resulting in suboptimal scene representation. Hence, reconstructing HR images from low-resolution (LR) observations is crucial for improving both human perception and subsequent applications.
@@
In contrast to upgrading hardware maintenance, super-resolution (SR) techniques provide a flexible and cost-effective alternative by predicting latent HR images from their LR counterparts [9, 10, 11, 12]. Early efforts often relied on hand-crafted priors to tame the ill-posedness [13, 14, 15, 16]. However, they struggled to produce accurate results and involved laborious optimization processes. Recently, deep neural networks have demonstrated remarkable progress in SR tasks and achieved superior performance over traditional approaches, such as Convolutional Neural Networks (CNNs) and Transformers. While CNN-based methods commonly invent elaborate attention mechanisms to grasp the informative features, restrained by the inherent nature of convolution units, they have limited respective fields and cannot capture long-range dependencies. As shown in Fig. 1, the effective receptive field (ERF) [17] of CNN-based model NLSN [18] is limited. This requirement is essential for SR tasks, as a predicted pixel needs prior knowledge from its surrounding region to be super-resolved, especially in wide-range RSI.
@@
Figure 1: The Effective Receptive Field (ERF) [17] comparison for (a) CNN-based method NLSN [18], (b) Transformer-based model RGT [19], and the proposed Mamba-based network FMSR. A wider distribution of dark areas demonstrates larger ERF. Our FMSR effectively obtains the largest ERF, indicating favorable global exploration capability.
@@
Transformer-based methods achieve increased respective fields by leveraging global interaction among all input data through a self-attention (SA) mechanism [20, 21, 22, 23, 24]. Despite achieving superior performance against CNN-based approaches, these methods exhibit quadratic complexity with respect to the token size. In this context, taming Transformer for high-resolution scenarios presents a significant challenge, particularly for large-scale remote sensing images. Although some approaches seek a lightweight SA for global modeling, such as recursive SA [19] and window-based SA [25], they usually come at the expense of global modeling accuracy and require stacking many blocks to establish a global dependency, thus increasing the computational budget. Moreover, the inherent issue of quadratic complexity remains unsolved. Therefore, a natural question arises: can a more efficient yet effective solution be developed to grasp the long-range dependencies across large-scale RSI?
@@
The recent-popular State Space Model (SSM) could be a promising answer to this question. Originating from Kalman filtering [26], SSM primarily employs linear filtering and prediction methods to represent the evolution of the internal state of the system, thus naturally enjoying linear complexity. By integrating SSM with the MLP block of Transformer, the simplified architecture Mamba [27] is achieved, which introduces a selection scanning mechanism in SSM to filter out irrelevant information for long-sequence modeling. Recently, Mamba demonstrated impressive results in various domains, making it a possible replacement for the Transformer model. Nevertheless, some potential problems persist in large-scale earth observation scenarios, making Mamba-based SR more challenging. Firstly, images captured by satellite platforms often lose crucial frequency information for perception, which requires heterogeneous representation for accurate reconstruction. The reason lies in that the original Mamba processes each token equally in the spatial domain, limiting its ability to perceive informative frequency signals. Secondly, there exists spatial diversity among observed objects in RSI, but is barely explored. Although the widely used channel attention mechanism could explore inter-channel relationships, it lacks explicit consideration of spatially-varying contents, resulting in suboptimal pixel-wise representation during local modeling.
@@
    We pioneer the first state space model for remote sensing image super-resolution (FMSR), demonstrating the potential of Mamba for efficient yet effective global modeling in large-scale remote sensing scenes.
@@
    To grasp the frequency dependencies, we devise a Frequency Selection Module (FSM) capable of adaptively selecting the most informative frequency cues through lightweight convolution and activation operations during Fourier transformation.
@@
    We develop a Hybrid Gate Module (HGM) that benefits the local bias of the CNN operator and spatially-varying coordinates to leverage more faithful localities for reconstruction.
@@ II-A Remote Sensing Image Super-resolution
RSI SR has witnessed significant advancements with the booming of deep learning [28]. The primary focus of this task lies in extracting prior knowledge from LR images, which can be broadly categorized into three categories: CNN-, Transformer-, and Mamba-based methods.
@@
Transformer-based. The core insight of Transformer lies in the Self-Attention (SA) mechanism, which has demonstrated superior long-range modeling capability and outperformed CNN-based methods. Lei et al. [35] devised a multi-stage Transformer-enhanced network. Recently, Chen et al. [25] proposed to activate more pixels in SR tasks by combining both CNN and self-attention. More Recently, an improved network [36] was proposed by integrating channel attention, improved SA, and anchored stripe attention. However, due to the quadratic complexity of SA regarding token size, they are less efficient in handling high-resolution images, large-scale RSI in particular. To alleviate the computational budget of SA, Chen et al. [19] proposed a recursive SA by recursively aggregating input features for enriched token representation. Nevertheless, efficient SA often sacrifices global modeling capability, and despite efforts to mitigate the quadratic complexity, the inherent problem of SA remains unsolved.
@@ III-B Frequency-assisted Mamba Block
$y={\alpha_{l}}\cdot{x_{l-1}}+{\rm VSSM}({\rm LN}({x_{l-1}}))+{\rm FAB}({x_{l-1}}).$ (5)
After that, $y$ undergoes further local modeling. Similarly, we adopt LN to normalize $y$ and then use a Hybrid Gate Module (HGM) to grasp the spatial locality. Also, FAB is employed to assist the local modeling process with frequency learning. Finally, another scale factor $\alpha_{l+1}$ is used to adaptively integrate the output from local and global representations, which can be expressed as:
$x_{l}={\beta_{l}}\cdot y+{\rm HGM}({\rm LN}(y))+{\rm FAB}(y).$ (6)
@@ III-C Vision State Space Module
Figure 5: Feature visualization comparisons. The feature maps corresponding to each reference image are the results of the 56th channels in the final FMG.
@@ III-D Hybrid Gate Module
Current methods often utilize either MLP layers [25] for feature propagation or incorporate convolution operations, such as attention mechanisms [37], after long-range exploration to introduce critical locality for improved performance. draws inspiration from these approaches but condenses them into a unified hybrid module. To learn a comprehensive local context, the HGM incorporates the spatially-varying properties of RSI by selective amplification or attenuation of local features on the pixel domain while preserving channel-specific features.
As shown in Fig. 3, HGM treats features captured by local convolution as coordinates and then multiplies them with a pixel-wise gating mask of the same size. Specifically, the input feature $\rm X$ is first fed through a $1\times 1$ convolution to expand the channel dimension to $2c$, then split by half in the channel dimension. In the first branch, a $1\times 1$ convolution, a $3\times 3$ depth-wise convolution, and a channel attention block [32] are used to yield the coordinates.
@@ III-E Frequency Selection Module
PSNR (dB) 27.691 27.985 27.178 28.154
@@
FMSR (Ours) - 30.93 0.8156 0.3001 34.04 0.8710 0.2325 30.97 0.8187 0.2983 31.98 0.8351 0.2770
FMSR++ (Ours) - 31.07 0.8185 0.3067 34.27 0.8735 0.2363 31.13 0.8219 0.3039 32.16 0.8380 0.2823
@@
Scene Categoties Bicubic EDSR [53] HSENet [34] NLSN [18] HAT-L [25] MambaIR [37] FMSR (Ours)
@@
Airport 27.83 0.7554 29.93 0.8282 30.08 0.8303 30.16 0.8322 30.15 0.8319 30.21 0.8327 30.26 0.8346
Bare Land 35.60 0.8564 36.94 0.8837 36.79 0.8841 37.00 0.8845 36.88 0.8841 36.95 0.8843 37.02 0.8847
Baseball Field 31.00 0.8305 33.05 0.8765 33.15 0.8774 33.24 0.8787 33.25 0.8789 33.32 0.8792 33.43 0.8803
Beach 32.90 0.8446 34.18 0.8727 34.14 0.8746 34.31 0.8749 34.34 0.8756 34.36 0.8754 34.38 0.8759
Bridge 30.22 0.8283 32.93 0.8800 33.06 0.8809 33.12 0.8818 33.04 0.8809 33.14 0.8821 33.28 0.8836
Center 26.51 0.6944 28.77 0.7921 28.83 0.7937 28.95 0.7971 28.92 0.7956 28.99 0.7966 29.09 0.8010
Church 24.29 0.6333 26.30 0.7469 26.47 0.7507 26.51 0.7528 26.56 0.7532 26.61 0.7543 26.64 0.7567
Commercial 27.33 0.7174 29.01 0.7940 29.21 0.7989 29.21 0.7996 29.21 0.8007 29.29 0.8012 29.34 0.8039
Dense Residential 22.93 0.5671 24.38 0.6839 24.60 0.6912 24.60 0.6936 24.67 0.6936 24.74 0.6965 24.78 0.7019
Desert 39.26 0.9100 40.20 0.9268 39.57 0.9271 40.27 0.9278 40.37 0.9278 40.09 0.9270 40.29 0.9275
Farmland 33.10 0.8226 35.00 0.8683 35.02 0.8692 35.10 0.8699 35.03 0.8691 35.09 0.8696 35.18 0.8711
Forest 28.79 0.6605 29.85 0.7315 30.00 0.7363 29.98 0.7369 30.01 0.7363 30.00 0.7370 30.04 0.7391
Industrial 26.77 0.6952 28.88 0.7931 28.98 0.7956 29.04 0.7982 29.04 0.7980 29.09 0.7988 29.17 0.8027
Meadow 33.86 0.7483 34.63 0.7804 34.55 0.7804 34.69 0.7821 34.70 0.7815 34.64 0.7811 34.74 0.7827
Medium Residential 26.36 0.6335 28.34 0.7365 28.45 0.7390 28.49 0.7418 28.46 0.7408 28.59 0.7435 28.65 0.7472
Mountain 29.51 0.7349 30.63 0.7885 30.72 0.7907 30.74 0.7916 30.78 0.7923 30.76 0.7918 30.78 0.7928
Park 29.06 0.7530 30.54 0.8130 30.71 0.8167 30.71 0.8177 30.71 0.8189 30.76 0.8185 30.80 0.8204
Parking 24.24 0.7060 27.25 0.8317 27.32 0.8341 27.57 0.8408 27.56 0.8405 27.65 0.8414 28.02 0.8524
Playground 32.64 0.8450 35.37 0.8943 35.46 0.8952 35.58 0.8967 35.49 0.8959 35.54 0.8961 35.72 0.8978
Pond 30.70 0.8167 32.11 0.8542 32.17 0.8549 32.22 0.8559 32.18 0.8555 32.24 0.8559 32.23 0.8560
Port 26.67 0.7986 28.50 0.8596 28.71 0.8623 28.71 0.8631 28.81 0.8638 28.83 0.8644 28.92 0.8668
Railway Station 26.78 0.6793 28.72 0.7738 28.84 0.7762 28.89 0.7783 28.88 0.7780 28.99 0.7802 29.04 0.7830
Resort 26.79 0.7029 28.52 0.7799 28.64 0.7825 28.68 0.7845 28.71 0.7849 28.76 0.7857 28.82 0.7886
River 30.37 0.7402 31.55 0.7891 31.61 0.7904 31.64 0.7914 31.63 0.7909 31.64 0.7912 31.66 0.7918
School 27.41 0.7237 29.36 0.8044 29.51 0.8074 29.55 0.8097 29.54 0.8104 29.62 0.8113 29.70 0.8143
Sparse Residential 26.66 0.6006 27.71 0.6728 27.84 0.6754 27.84 0.6767 27.88 0.6759 27.88 0.6768 27.91 0.6789
Square 28.55 0.7391 30.84 0.8200 30.94 0.8223 31.04 0.8244 31.00 0.8251 31.05 0.8247 31.16 0.8276
Stadium 27.16 0.7547 29.63 0.8387 29.68 0.8391 29.79 0.8422 29.77 0.8422 29.83 0.8421 29.97 0.8460
Storage Tanks 25.65 0.6793 27.44 0.7664 27.58 0.7688 27.61 0.7709 27.60 0.7698 27.64 0.7709 27.71 0.7739
Viaduct 26.97 0.6755 28.99 0.7757 29.08 0.7772 29.17 0.7813 29.11 0.7794 29.19 0.7810 29.26 0.7840
Average 28.86 0.7382 30.65 0.8086 30.72 0.8108 30.81 0.8126 30.81 0.8124 30.85 0.8130 30.93 0.8156

### v2

@@ Abstract
Recent progress in remote sensing image (RSI) super-resolution (SR) has exhibited remarkable performance using deep neural networks, e.g., Convolutional Neural Networks and Transformers. However, existing SR methods often suffer from either a limited receptive field or quadratic computational overhead, resulting in sub-optimal global representation and unacceptable computational costs in large-scale RSI. To alleviate these issues, we develop the first attempt to integrate the Vision State Space Model (Mamba) for RSI-SR, which specializes in processing large-scale RSI by capturing long-range dependency with linear complexity. To achieve better SR reconstruction, building upon Mamba, we devise a Frequency-assisted Mamba framework, dubbed FMSR, to explore the spatial and frequent correlations. In particular, our FMSR features a multi-level fusion architecture equipped with the Frequency Selection Module (FSM), Vision State Space Module (VSSM), and Hybrid Gate Module (HGM) to grasp their merits for effective spatial-frequency fusion. Considering that global and local dependencies are complementary and both beneficial for SR, we further recalibrate these multi-level features for accurate feature fusion via learnable scaling adaptors. Extensive experiments on AID, DOTA, and DIOR benchmarks demonstrate that our FMSR outperforms state-of-the-art Transformer-based methods HAT-L in terms of PSNR by 0.11 dB on average, while consuming only 28.05% and 19.08% of its memory consumption and complexity, respectively. Code will be available at https://github.com/XY-boy/FreMamba
@@ I Introduction
High-resolution remote sensing imagery (RSI), which records high-quality earth observation details, provides promising prospects for large-scale and fine-grained applications [1, 2, 3, 4, 5, 6, 7]. However, the complex imaging environment (e.g., scattering and tremor) often impedes high-resolution (HR) image acquisition [8, 9, 10, 11]. Moreover, RSI often undergoes severe compression and downsampling to tame the transmission instability between satellites and ground stations, resulting in suboptimal scene representation. Hence, reconstructing HR images from low-resolution (LR) observations is crucial for improving both human perception and subsequent applications.
@@
In contrast to upgrading hardware maintenance, super-resolution (SR) techniques provide a flexible and cost-effective alternative by predicting latent HR images from their LR counterparts [12, 13, 14, 15]. Early efforts often relied on hand-crafted priors to tame the ill-posedness [16, 17, 18, 19, 20]. However, they struggled to produce accurate results and involved laborious optimization processes. Recently, deep neural networks have demonstrated remarkable progress in SR tasks and achieved superior performance over traditional approaches, such as Convolutional Neural Networks (CNNs) and Transformers. While CNN-based methods commonly invent elaborate attention mechanisms to grasp the informative features, restrained by the inherent nature of convolution units, they have limited respective fields and cannot capture long-range dependencies. As shown in Fig. 1, the effective receptive field (ERF) [21] of CNN-based model NLSN [22] is limited. This requirement is essential for SR tasks, as a predicted pixel needs prior knowledge from its surrounding region to be super-resolved, especially in wide-range RSI.
@@
Figure 1: The Effective Receptive Field (ERF) [21] comparison for (a) CNN-based method NLSN [22], (b) Transformer-based model ATD [23], and the proposed Mamba-based network FMSR. A wider distribution of dark areas demonstrates larger ERF. Our FMSR effectively obtains the largest ERF, indicating favorable global exploration capability.
@@
Transformer-based methods achieve increased respective fields by leveraging global interaction among all input data through a self-attention (SA) mechanism, demonstrating impressive performance across various domains [24, 25, 26, 27, 28, 29]. Despite achieving superior performance against CNN-based approaches, these methods exhibit quadratic complexity with respect to the token size. In this context, taming Transformer for high-resolution scenarios presents a significant challenge, particularly for large-scale remote sensing images. Although some approaches seek a lightweight SA for global modeling, such as recursive SA [30] and window-based SA [31], they usually come at the expense of global modeling accuracy and require stacking many blocks to establish a global dependency, thus increasing the computational budget. Moreover, the inherent issue of quadratic complexity remains unsolved. Therefore, a natural question arises: can a more efficient yet effective solution be developed to grasp the long-range dependencies across large-scale RSI?
@@
The recent-popular State Space Model (SSM) could be a promising answer to this question. Originating from Kalman filtering [32], SSM primarily employs linear filtering and prediction methods to represent the evolution of the internal state of the system, thus naturally enjoying linear complexity. By integrating SSM with the MultiLayer Perceptron (MLP) block of Transformer, the simplified architecture Mamba [33] is achieved, which introduces a selection scanning mechanism in SSM to filter out irrelevant information for long-sequence modeling. Recently, Mamba demonstrated impressive results in various domains, making it a possible replacement for the Transformer model. Nevertheless, although Mamba exhibits favorable performance and can serve as an alternative to Transformer, some potential problems persist in large-scale earth observation scenarios, making introducing Mamba into RSI SR more challenging. Firstly, images captured by satellite platforms often lose crucial frequency information for perception, which requires heterogeneous representation for accurate reconstruction. The underlying reason lies in that the original Mamba processes each token equally in the spatial domain, limiting its ability to perceive informative frequency signals across entire images. Secondly, there exists spatial diversity among observed objects in RSI, which is barely explored. While Mamba is effective in exploring long-range dependencies, it lacks explicit consideration of spatially-varying contents, resulting in suboptimal pixel-wise representation during local modeling.
@@
    We introduce the first state space model for remote sensing image super-resolution (FMSR), highlighting Mamba’s capability for efficient and effective global modeling in large-scale remote sensing scenarios.
@@
    To integrate more high-frequency cues into Mamba, we develop a Frequency Selection Module (FSM), which adaptively identifies and selects the most informative frequency signals during the fast fourier transformation process.
@@
    We design a Hybrid Gate Module (HGM) that integrates the local bias of CNN operators with spatially-varying coordinates to enhance the locality of feature representation, leading to more accurate and faithful SR performance.
@@ II-A Remote Sensing Image Super-resolution
RSI SR has witnessed significant advancements with the booming of deep learning [34, 35, 36, 37, 38]. The primary focus of this task lies in extracting prior knowledge from LR images, which can be broadly categorized into three categories: CNN-, Transformer-, and Mamba-based methods.
@@
Transformer-based. The core insight of Transformer lies in the Self-Attention (SA) mechanism [45], which has demonstrated superior long-range modeling capability and outperformed CNN-based methods. Lei et al. [46] devised a multi-stage Transformer-enhanced network. Recently, Chen et al. [31] proposed to activate more pixels in SR tasks by combining both CNN and self-attention. More Recently, an improved network [47] was proposed by integrating channel attention, improved SA, and anchored stripe attention. However, due to the quadratic complexity of SA regarding token size, they are less efficient in handling high-resolution images, large-scale RSI in particular. To alleviate the computational budget of SA, Chen et al. [30] proposed a recursive SA by recursively aggregating input features for enriched token representation. Nevertheless, efficient SA often sacrifices global modeling capability, and despite efforts to mitigate the quadratic complexity, the inherent problem of SA remains unsolved.
@@ III-B Frequency-assisted Mamba Block
$y={\alpha_{l}}\cdot{x_{l-1}}+{\rm VSSM}({\rm LN}({x_{l-1}}))+{\rm FSM}({x_{l-1}}).$ (5)
After that, $y$ undergoes further local modeling. Similarly, we adopt LN to normalize $y$ and then use a Hybrid Gate Module (HGM) to grasp the spatial locality. Also, FSM is employed to assist the local modeling process with frequency learning. Finally, another scale factor $\alpha_{l+1}$ is used to adaptively integrate the output from local and global representations, which can be expressed as:
$x_{l}={\alpha_{l+1}}\cdot y+{\rm HGM}({\rm LN}(y))+{\rm FSM}(y).$ (6)
@@ III-C Vision State Space Module
Figure 5: Feature visualization comparisons. The feature maps corresponding to each reference image are the results of the 56-th channels in the final FMG.
@@ III-D Hybrid Gate Module
Current methods often utilize either MLP layers [31] for feature propagation or incorporate convolution operations, such as attention mechanisms [48], after long-range exploration to introduce critical locality for improved performance. Our approach draws inspiration from these methods but condenses them into a unified hybrid module. To learn a comprehensive local context, the Hybrid Gate Module (HGM) incorporates the spatially-varying properties of remote sensing images (RSI) by selectively amplifying or attenuating local features in the pixel domain while preserving channel-specific features.
As shown in Fig. 3, HGM treats features captured by local convolution as coordinates and then multiplies them with a pixel-wise gating mask of the same size. Specifically, the input feature $\rm X$ is first processed through a $1\times 1$ convolution to expand the channel dimension to $2c$. We then split $\rm X$ into two parts, $\rm X_{1}$ and $\rm X_{2}$, by halving the channel dimension. $\rm X_{1}$ and $\rm X_{2}$ are subsequently sent into the first and second branches, respectively. In the first branch, a $1\times 1$ convolution, a $3\times 3$ depth-wise convolution, and a channel attention block [42] are used to yield the coordinates.
@@ III-E Frequency Selection Module
PSNR (dB) 27.691 27.985 28.178 28.154
Figure 7: Ablation studies on the effect of different expansion factors $\lambda$.
Figure 8: Computational complexity comparison with inputs of different resolutions. We adopt the standard Multihead Self-Attention [1] as the baseline. Initially, we adjust the model to ensure that the consumption of parameters and FLOPs is roughly equivalent. Subsequently, we increase the input resolution from $32\times 32$ to $88\times 8$ 88.
@@
FMSR (Ours) TMM’2024 30.93 0.8156 0.3001 34.04 0.8710 0.2325 30.97 0.8187 0.2983 31.98 0.8351 0.2770
FMSR++ (Ours) TMM’2024 31.07 0.8185 0.3067 34.27 0.8735 0.2363 31.13 0.8219 0.3039 32.16 0.8380 0.2823
@@
Categoties Bicubic EDSR [64] HSENet [44] NLSN [22] HAT-L [31] MambaIR [48] RGT [30] FMSR (Ours)
@@
Airport 27.83 0.7554 29.93 0.8282 30.08 0.8303 30.16 0.8322 30.15 0.8319 30.21 0.8327 30.29 0.8352 30.26 0.8346
Bare Land 35.60 0.8564 36.94 0.8837 36.79 0.8841 37.00 0.8845 36.88 0.8841 36.95 0.8843 37.00 0.8852 37.02 0.8847
Baseball Field 31.00 0.8305 33.05 0.8765 33.15 0.8774 33.24 0.8787 33.25 0.8789 33.32 0.8792 33.35 0.8805 33.43 0.8803
Beach 32.90 0.8446 34.18 0.8727 34.14 0.8746 34.31 0.8749 34.34 0.8756 34.36 0.8754 34.41 0.8766 34.38 0.8759
Bridge 30.22 0.8283 32.93 0.8800 33.06 0.8809 33.12 0.8818 33.04 0.8809 33.14 0.8821 33.22 0.8832 33.28 0.8836
Center 26.51 0.6944 28.77 0.7921 28.83 0.7937 28.95 0.7971 28.92 0.7956 28.99 0.7966 29.09 0.8015 29.09 0.8010
Church 24.29 0.6333 26.30 0.7469 26.47 0.7507 26.51 0.7528 26.56 0.7532 26.61 0.7543 26.66 0.7580 26.64 0.7567
Commercial 27.33 0.7174 29.01 0.7940 29.21 0.7989 29.21 0.7996 29.21 0.8007 29.29 0.8012 29.30 0.8048 29.34 0.8039
D-Residential 22.93 0.5671 24.38 0.6839 24.60 0.6912 24.60 0.6936 24.67 0.6936 24.74 0.6965 24.78 0.7019 24.78 0.7019
Desert 39.26 0.9100 40.20 0.9268 39.57 0.9271 40.27 0.9278 40.37 0.9278 40.09 0.9270 40.15 0.9283 40.29 0.9275
Farmland 33.10 0.8226 35.00 0.8683 35.02 0.8692 35.10 0.8699 35.03 0.8691 35.09 0.8696 35.14 0.8715 35.18 0.8711
Forest 28.79 0.6605 29.85 0.7315 30.00 0.7363 29.98 0.7369 30.01 0.7363 30.00 0.7370 30.01 0.7400 30.04 0.7391
Industrial 26.77 0.6952 28.88 0.7931 28.98 0.7956 29.04 0.7982 29.04 0.7980 29.09 0.7988 29.20 0.8034 29.17 0.8027
Meadow 33.86 0.7483 34.63 0.7804 34.55 0.7804 34.69 0.7821 34.70 0.7815 34.64 0.7811 34.69 0.7831 34.74 0.7827
M-Residential 26.36 0.6335 28.34 0.7365 28.45 0.7390 28.49 0.7418 28.46 0.7408 28.59 0.7435 28.66 0.7479 28.65 0.7472
Mountain 29.51 0.7349 30.63 0.7885 30.72 0.7907 30.74 0.7916 30.78 0.7923 30.76 0.7918 30.71 0.7941 30.78 0.7928
Park 29.06 0.7530 30.54 0.8130 30.71 0.8167 30.71 0.8177 30.71 0.8189 30.76 0.8185 30.81 0.8211 30.80 0.8204
Parking 24.24 0.7060 27.25 0.8317 27.32 0.8341 27.57 0.8408 27.56 0.8405 27.65 0.8414 27.79 0.8472 28.02 0.8524
Playground 32.64 0.8450 35.37 0.8943 35.46 0.8952 35.58 0.8967 35.49 0.8959 35.54 0.8961 35.65 0.8978 35.72 0.8978
Pond 30.70 0.8167 32.11 0.8542 32.17 0.8549 32.22 0.8559 32.18 0.8555 32.24 0.8559 32.24 0.8567 32.23 0.8560
Port 26.67 0.7986 28.50 0.8596 28.71 0.8623 28.71 0.8631 28.81 0.8638 28.83 0.8644 28.90 0.8667 28.92 0.8668
Railway Station 26.78 0.6793 28.72 0.7738 28.84 0.7762 28.89 0.7783 28.88 0.7780 28.99 0.7802 29.03 0.7833 29.04 0.7830
Resort 26.79 0.7029 28.52 0.7799 28.64 0.7825 28.68 0.7845 28.71 0.7849 28.76 0.7857 28.80 0.7889 28.82 0.7886
River 30.37 0.7402 31.55 0.7891 31.61 0.7904 31.64 0.7914 31.63 0.7909 31.64 0.7912 31.67 0.7927 31.66 0.7918
School 27.41 0.7237 29.36 0.8044 29.51 0.8074 29.55 0.8097 29.54 0.8104 29.62 0.8113 29.71 0.8150 29.70 0.8143
S-Residential 26.66 0.6006 27.71 0.6728 27.84 0.6754 27.84 0.6767 27.88 0.6759 27.88 0.6768 27.91 0.6796 27.91 0.6789
Square 28.55 0.7391 30.84 0.8200 30.94 0.8223 31.04 0.8244 31.00 0.8251 31.05 0.8247 31.16 0.8284 31.16 0.8276
Stadium 27.16 0.7547 29.63 0.8387 29.68 0.8391 29.79 0.8422 29.77 0.8422 29.83 0.8421 29.94 0.8457 29.97 0.8460
Storage Tanks 25.65 0.6793 27.44 0.7664 27.58 0.7688 27.61 0.7709 27.60 0.7698 27.64 0.7709 27.71 0.7744 27.71 0.7739
Viaduct 26.97 0.6755 28.99 0.7757 29.08 0.7772 29.17 0.7813 29.11 0.7794 29.19 0.7810 29.26 0.7851 29.26 0.7840
Average 28.86 0.7382 30.65 0.8086 30.72 0.8108 30.81 0.8126 30.81 0.8124 30.85 0.8130 30.91 0.8159 30.93 0.8156